# Barrow holographic dark energy with varying exponent

### v1

@@ 1 Introduction
An alternative explanation for the origin of dark energy emerges through the cosmological application of the holographic principle [8, 9, 10]. This framework, rooted in the thermodynamics of black holes, connects the ultraviolet cutoff of a quantum field theory with the largest distance of the theory, a prerequisite for its applicability at large distances [11]. In a region where entropy is proportional to volume, the total energy should not exceed the mass of a black hole with the same radius, and this saturation leads to the extraction of a holographically originated vacuum energy, and thus to a form of holographic dark energy with a dynamic nature [12, 13]. The cosmological implications of holographic dark energy are intriguing [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] and align with observations [23, 24, 25, 26, 27, 28]. Additionally, one has the advantage that holographic scenarios are free from potential pathologies that may appear in modified gravity [7]. Hence, holographic dark energy has spurred extensive research, leading to various extensions [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58].
@@
All the above extended-entropy scenarios incorporate a constant parameter that quantifies the deviation form standard entropy. However, in principle one expects a dynamical, energy-scale-dependent behavior for this parameter, aligning with the typical case observed in quantum field theory and quantum gravity under renormalization group applications. In particular, all parameters and coupling constants demonstrate a running nature in tandem with the energy scale, and in a cosmological framework this would effectively generate a time-dependence. A first investigation of holographic dark energy models with extended entropy with varying exponent was performed in [73] and it was shown that the running behavior can lead to interesting physical implications.
From the aforementioned extended entropies, Barrow entropy has been shown to lead to a very interesting cosmological phenomenology. It arises from quantum-gravitational phenomena on the horizon structure, parametrized by the single exponent $\Delta$ [60], where $\Delta=0$ corresponds to standard entropy and $\Delta=1$ to maximal deviation. Hence, it is natural to consider that at early times these quantum phenomena are more intense and thus $\Delta$ is closer to 1, while as time passes they smooth out and $\Delta$ tends to its standard value 0. In summary, a scenario of Barrow entropy with varying exponent has even more justification than the usual energy-scale dependence of coupling constants.
In the present work we are interested in studying Barrow holographic dark energy with varying exponent. In particular, we impose various parametrizations for the time-dependence, which is equivalent to redshift-dependence, and we examine their effect on the cosmological evolution. The plan of the work is the following. In Section 2 we review standard Barrow holographic dark energy, and then in Section 3 we construct the extended scenario where the Barrow exponent is varying. In Section 4 we impose specific ansätze for the Barrow running exponent and we investigate their cosmological behavior. Finally, in Section 5 we summarize our results.
@@ 3 Barrow holographic dark energy with varying exponent
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{\Omega_{DE}^{\prime}}{\Omega_{DE}\left(1-\Omega_{DE}\right)}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\Delta}{2(\Delta-2)}}(2-\Delta)+\Delta+1$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{\Delta^{{}^{\prime}}}{\Delta-2}},$ (3.1)
@@
$Q=Q(x)\equiv\frac{Ce^{3x}}{3M_{p}^{2}H_{0}^{2}\Omega_{m0}}.$ (3.2)
@@
$w_{DE}=-\frac{\Delta\!+\!1}{3}+\frac{(\Delta\!-\!2)\sqrt{\Omega_{DE}}}{3}\left(\!\frac{{C}}{3M_{p}^{2}}\!\right)^{\!-\frac{1}{2}}\!\left[\frac{Q(1\!-\!\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\Delta}{2(\Delta\!-\!2)}}-\frac{\Delta^{{}^{\prime}}}{3}\log{\left[\frac{Q(1\!-\!\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2\!-\!\Delta}}}.$ (3.4)
@@ 4.1 Linear case
$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta z}{2(\alpha+\beta z-2)}}(2-\alpha-\beta z)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{-\beta(1+z)}{\alpha+\beta z-2}}+\alpha+\beta z+1,\hskip 28.00006pt$ (4.1)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta z+1}{3}+\frac{(\alpha+\beta z-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta z}{2(\alpha+\beta z-2)}}$
$\ \ \ +\frac{(1+z)\beta}{3}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta z}}}.$ (4.2)
@@ 4.2 CPL-like case
$\!\!\!\!\!\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha(z+1)+\beta z}{2(\alpha-2)(z+1)+2\beta z}}\left[2\!-\!\alpha\!-\!\frac{\beta z}{z+1}\right]$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ \ \ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{-\frac{\beta}{(\alpha-2)(z+1)+\beta z}}+\alpha+\frac{\beta z}{z+1}+1,\hskip 28.00006pt$ (4.4)
@@
$\!\!\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\frac{\beta z}{z+1}+1}{3}+\frac{\left(\alpha+\frac{\beta z}{z+1}-2\right)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha(z+1)+\beta z}{2(\alpha-2)(z+1)+2\beta z}}$
$\ \ \ \ \ -\frac{\beta}{3(z+1)}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{z+1}{(2-\alpha)(z+1)-\beta z}}}.$ (4.5)
@@ 4.3 Exponential case
$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta e^{-\lambda z}}{2(\alpha+\beta e^{-\lambda z}-2)}}[2-(\alpha+\beta e^{-\lambda z})]$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{(1+z)\beta\lambda e^{-\lambda z}}{\alpha+\beta e^{-\lambda z}-2}}+\alpha+\beta e^{-\lambda z}+1,\hskip 28.00006pt$ (4.6)
@@
$\!\!\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta e^{-\lambda z}+1}{3}+\frac{(\alpha+\beta e^{-\lambda z}-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta e^{-\lambda z}}{2(\alpha+\beta e^{-\lambda z}-2)}}$
$\ \ \ \ \ -\frac{(1+z)\beta\lambda e^{-\lambda z}}{3}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta e^{-\lambda z}}}}.$ (4.7)
@@ 4.4 Hyperbolic Tangent case
$\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\tanh({\gamma z})}{2[\alpha+\beta\tanh({\gamma z})-2)]}}[2-\alpha-\beta\tanh({\gamma z})]$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{-(1+z)\beta\gamma\text{sech}^{2}({\gamma z})}{\alpha+\beta\tanh({\gamma z})-2}}+\alpha+\beta\tanh({\gamma z})+1,\hskip 28.00006pt$ (4.8)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!w_{DE}=\frac{[\alpha\!+\!\beta\tanh({\gamma z})\!-\!2]\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\!\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\tanh({\gamma z})}{2[\alpha\!+\!\beta\tanh({\gamma z})-2)]}}$
$\ \ \ \,-\frac{\alpha\!+\!\beta\tanh({\gamma z})\!+\!1}{3}+\frac{(1+z)\beta\gamma\,\text{sech}^{2}({\gamma z})}{3}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2\!-\!\alpha\!-\!\beta\tanh{\gamma z}}}}.$ (4.9)

### v2

@@ 1 Introduction
An alternative explanation for the origin of dark energy emerges through the cosmological application of the holographic principle [8, 9, 10]. This framework, rooted in the thermodynamics of black holes, connects the ultraviolet cutoff of a quantum field theory with the largest distance of the theory, a prerequisite for its applicability at large distances [11]. In a region where entropy is proportional to volume, the total energy should not exceed the mass of a black hole with the same radius, and this saturation leads to the extraction of a holographically originated vacuum energy, and thus to a form of holographic dark energy with a dynamic nature [12, 13]. The cosmological implications of holographic dark energy are intriguing [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22] and align with observations [23, 24, 25, 26, 27, 28]. Additionally, one has the advantage that holographic scenarios are free from potential pathologies that may appear in modified gravity [7]. Hence, holographic dark energy has spurred extensive research, leading to various extensions [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60].
@@
All the above extended-entropy scenarios incorporate a constant parameter that quantifies the deviation form standard entropy. However, in principle one expects a dynamical, energy-scale-dependent behavior for this parameter, aligning with the typical case observed in quantum field theory and quantum gravity under renormalization group applications. In particular, all parameters and coupling constants demonstrate a running nature in tandem with the energy scale, and in a cosmological framework this would effectively generate a time-dependence. A first investigation of Tsallis holographic dark energy models with extended entropy with varying exponent was performed in [75] and it was shown that the running behavior can lead to interesting physical implications.
Nevertheless, in the case of Barrow entropy one has an additional necessity for the varying behavior. Since it arises from quantum-gravitational phenomena on the horizon structure, parametrized by the single exponent $\Delta$ [62], where $\Delta=0$ corresponds to standard entropy and $\Delta=1$ to maximal deviation, it is hard to justify why in the recent Universe one has significant quantum-gravitational phenomena in its horizon that could lead to significant deviation from standard, i.e. classical entropy. Hence, it is more natural to consider that at early times these quantum phenomena are more intense and thus $\Delta$ is closer to 1, while as time passes they smooth out and $\Delta$ tends to its standard value 0. We mention here that such a scenario is still Barrow entropy, since $\Delta$ does arise from quantum gravitational phenomena, however one goes beyond the first approximation of [62] in which these phenomena are constant throughout the Universe evolution, and examines more realistic scenarios in which quantum gravitational phenomena are stronger at early times and smooth out as the system becomes larger.
In summary, a scenario of Barrow entropy with varying exponent has even more justification than the usual energy-scale dependence of coupling constants.
Since Barrow entropy has been shown to lead to a very interesting cosmological phenomenology, in the present work we are interested in studying Barrow holographic dark energy with varying exponent. In particular, we impose various parametrizations for the time-dependence, which is equivalent to redshift-dependence, and we examine their effect on the cosmological evolution. The plan of the work is the following. In Section 2 we review standard Barrow holographic dark energy, and then in Section 3 we construct the extended scenario where the Barrow exponent is varying. In Section 4 we impose specific ansätze for the Barrow running exponent and we investigate their cosmological behavior. Finally, in Section 5 we summarize our results.
@@ 3 Barrow holographic dark energy with varying exponent
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{\Omega_{DE}^{\prime}}{\Omega_{DE}\left(1-\Omega_{DE}\right)}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\Delta}{2(\Delta-2)}}(2-\Delta)+\Delta+1$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{\Delta^{{}^{\prime}}}{\Delta-2}},$ (3.1)
@@
$P=P(x)\equiv\frac{Ce^{3x}}{3M_{p}^{2}H_{0}^{2}\Omega_{m0}}.$ (3.2)
@@
$w_{DE}=-\frac{\Delta\!+\!1}{3}+\frac{(\Delta\!-\!2)\sqrt{\Omega_{DE}}}{3}\left(\!\frac{{C}}{3M_{p}^{2}}\!\right)^{\!-\frac{1}{2}}\!\left[\frac{P(1\!-\!\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\Delta}{2(\Delta\!-\!2)}}-\frac{\Delta^{{}^{\prime}}}{3}\log{\left[\frac{P(1\!-\!\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2\!-\!\Delta}}}.$ (3.4)
@@ 4 Cosmological evolution
We mention here that we parameterize the way that quantum gravitational phenomena smooth out, i.e $\Delta$ as a function of redshift, in a purely phenomenological manner, through various functions. Definitely each of these functions at the fundamental level should correspond to a particular way that quantum gravitational phenomena behave.
@@ 4.1 Linear case
$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta z}{2(\alpha+\beta z-2)}}(2-\alpha-\beta z)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{-\beta(1+z)}{\alpha+\beta z-2}}+\alpha+\beta z+1,\hskip 28.00006pt$ (4.1)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta z+1}{3}+\frac{(\alpha+\beta z-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta z}{2(\alpha+\beta z-2)}}$
$\ \ \ +\frac{(1+z)\beta}{3}\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta z}}}.$ (4.2)
@@ 4.2 CPL-like case
$\!\!\!\!\!\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha(z+1)+\beta z}{2(\alpha-2)(z+1)+2\beta z}}\left[2\!-\!\alpha\!-\!\frac{\beta z}{z+1}\right]$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ \ \ \ \ \ \ +\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{-\frac{\beta}{(\alpha-2)(z+1)+\beta z}}+\alpha+\frac{\beta z}{z+1}+1,\hskip 28.00006pt$ (4.4)
@@
$\!\!\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\frac{\beta z}{z+1}+1}{3}+\frac{\left(\alpha+\frac{\beta z}{z+1}-2\right)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha(z+1)+\beta z}{2(\alpha-2)(z+1)+2\beta z}}$
$\ \ \ \ \ -\frac{\beta}{3(z+1)}\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{z+1}{(2-\alpha)(z+1)-\beta z}}}.$ (4.5)
@@ 4.3 Exponential case
$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta e^{-\lambda z}}{2(\alpha+\beta e^{-\lambda z}-2)}}[2-(\alpha+\beta e^{-\lambda z})]$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ \ \ \ \ +\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{(1+z)\beta\lambda e^{-\lambda z}}{\alpha+\beta e^{-\lambda z}-2}}+\alpha+\beta e^{-\lambda z}+1,\hskip 28.00006pt$ (4.6)
@@
$\!\!\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta e^{-\lambda z}+1}{3}+\frac{(\alpha+\beta e^{-\lambda z}-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta e^{-\lambda z}}{2(\alpha+\beta e^{-\lambda z}-2)}}$
$\ \ \ \ \ -\frac{(1+z)\beta\lambda e^{-\lambda z}}{3}\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta e^{-\lambda z}}}}.$ (4.7)
@@ 4.4 Hyperbolic Tangent case
$\!\!\!\!\frac{-(1+z)}{\Omega_{DE}\left(1-\Omega_{DE}\right)}\frac{d\Omega_{DE}}{dz}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\tanh({\gamma z})}{2[\alpha+\beta\tanh({\gamma z})-2)]}}[2-\alpha-\beta\tanh({\gamma z})]$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{-(1+z)\beta\gamma\text{sech}^{2}({\gamma z})}{\alpha+\beta\tanh({\gamma z})-2}}+\alpha+\beta\tanh({\gamma z})+1,\hskip 28.00006pt$ (4.8)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!w_{DE}=\frac{[\alpha\!+\!\beta\tanh({\gamma z})\!-\!2]\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\!\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\tanh({\gamma z})}{2[\alpha\!+\!\beta\tanh({\gamma z})-2)]}}$
$\ \ \ \,-\frac{\alpha\!+\!\beta\tanh({\gamma z})\!+\!1}{3}+\frac{(1+z)\beta\gamma\,\text{sech}^{2}({\gamma z})}{3}\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2\!-\!\alpha\!-\!\beta\tanh{\gamma z}}}}.$ (4.9)
@@
We close this subsection by confronting the model of the hyperbolic tangent case at hand with Supernovae type Ia (SNIa) and Cosmic Chronometer data. For SNIa, it is well established that the relation between the apparent and absolute magnitudes follows
$2.5\log\left[\frac{L}{l(z)}\right]=\mu\equiv m(z)-M=5\log\left[\frac{d_{L}(z)_{\text{obs}}}{Mpc}\right]+25~{},$ (4.10)
with $l(z)$ and $m(z)$ the apparent luminosity and apparent magnitude, and $L$ and $M$ the absolute luminosity and magnitude, respectively, while $d_{L}(z)_{\text{obs}}$ is the luminosity distance. Moreover, the theoretical expression for the luminosity distance is given by
$d_{L}\left(z\right)_{\text{th}}\equiv\left(1+z\right)\int^{z}_{0}\frac{dz^{\prime}}{H\left(z^{\prime}\right)}~{}.$ (4.11)
Figure 5: The apparent minus absolute magnitude predicted theoretically in the Hyperbolic Tangent case (red - dashed) with $\alpha=\beta=1/2$ and with $\gamma=0.001$, in units where $M_{p}^{2}=1$, on top of the Pantheon SNIa data points from [96]. For comparison we depict the $\Lambda$CDM curve (black - solid) too.
Figure 6: The Hubble parameter $H(z)$ in units of Km/s/Mpc as a function of the redshift, for the hyperbolic tangent case with $\alpha=\beta=1/2$ and with $\gamma=0.001$ (red-solid), in $M_{p}^{2}=1$ units, on top of the Cosmic Chronometers data points from [97] at $2\sigma$ confidence level. For comparison we depict the $\Lambda$CDM curve (black - solid) too. We have imposed $\Omega_{m0}\approx 0.31$.
Since the hyperbolic tangent case, as well as the $\Lambda$CDM scenario, provide predictions for $H(z)$, we compare the theoretically computed apparent minus absolute magnitudes with the binned Pantheon SNIa data from [96] in Fig. 5. As illustrated, the agreement is very good, with our model accurately reproducing the observed cosmological behavior.
In addition, the Cosmic Chronometer (CC) dataset, based on the measurement of $H(z)$ from the relative ages of passively evolving galaxies, allows us to compare the observed $H(z)$ values with our model predictions. In Fig. 6, we show the comparison between the theoretical evolution of $H(z)$ in our model with a hyperbolic tangent form and the $\Lambda$CDM model with the CC data points from [97], presented at a $2\sigma$ confidence level. The agreement is again very good, with our model reproducing the observed accelerating expansion for the parameter set $\{\alpha,\beta,\gamma\}=\{0.5,0.5,0.001\}$.
In summary, there exist regions in the parameter space of our model that are able to reproduce the observed evolution of the Hubble function. This suggests that this model is a viable candidate. A more thorough evaluation, including likelihood analysis and model selection using complete cosmological datasets, will be presented in future work.